OpenStudy (anonymous):

What is a simpler form of the radical expression? 4^sqrt 81x^20y^8 B. 9|x^25|y^4 C.3|x^5|y^2 D. 9x^25|y^4|

OpenStudy (agent0smith):

Can you take a screenshot? Cos i have no idea...

OpenStudy (agent0smith):

\[\Large \sqrt[4]{81x^{20} y^8}\]

OpenStudy (agent0smith):

start by writing everything in terms of powers of 4 \[\Large \sqrt[4]{3^4 (x^{5})^4 (y^2)^4}\]

OpenStudy (muzzack):

\(\huge\ 3\sqrt[4]{x^{20}y^8}\)

OpenStudy (muzzack):

C is ur answer my dear

OpenStudy (agent0smith):

She wants to actually learn, not be given answers @Muzzack

OpenStudy (anonymous):

Yeah lol

OpenStudy (muzzack):

sorry sire

OpenStudy (muzzack):

ill leave this to you @agent0smith im terrible sorry

OpenStudy (anonymous):

Hahahahaa its okay cx

OpenStudy (anonymous):

Okay so now what?

OpenStudy (agent0smith):

\[\Large \sqrt[4]{3^4 (x^{5})^4 (y^2)^4}\]now you can simplify... just like a square root cancels out a square, a 4th root cancels out a power of 4.

OpenStudy (anonymous):

Oh okay so 3(x)(y^2) ?

OpenStudy (anonymous):

Or is that so wrong ?

OpenStudy (agent0smith):

Close but not quite. The only things that cancel are powers of 4

OpenStudy (anonymous):

Oh so its (x^5) right ?

OpenStudy (agent0smith):

Just write the whole expression as you did a minute ago

OpenStudy (anonymous):

Oh okay ummm 3(x^5)(y^2)?